Solve dosage problems using basic calculation, ratio and proportion methods. Dose ordered verses dose on hand calculations.
Answer #1

Ratio and Proportion method:

Ratio and proportion are one of the oldest methods in calculating medication dosage. In this, the known quantities are kept at left side and right with the desired dose and amount to be given.

H = Dose on hand

V = volume of the drug

D = Desired dose(ordered)

X = Amount to be given

Then, it can be displayed as

H : V D : X

Then multiply the extremes and means as follows

HX = DV

Then X (amount) can find out by

X = DV / H.

For example, Dose on hand is 500mg in 3mL. The doctors ordered for 750mg, then the amount of drug will be X

Dose on Hand H = 500mg

Volume of drug V = 3mL

Ordered dose D = 750mg

Amount to give X = ?

500 : 3 750 : X

By crossing the extremes and means

500X = 3 *750

500X = 2250

X = 2250 / 500

X = 4.5mL

Basic formula method

It can also be calculated using the basic formula

Desireddose(D)) x Quantity(V) = Amount Doseon Hand(H))

For example,

D = 750mg

H = 500mg

Q = 3mL

then,

750 500 * 3 = 4.5mL